# Learning large softmax mixtures with warm start EM

### v1

@@ 2.5 Practical choice of the primary axis
so that the selected $v_{i^{*}}$ approximately maximizes $\det({\bm{M}}(v_{i}))$, thereby leading to the largest separation among $v_{i}^{\top}\Sigma{\bm{\theta}}_{1}^{*},\ldots,v_{i}^{\top}\Sigma{\bm{\theta}}_{K}^{*}.$
We conclude this section by noting that the above rates are satisfactory in theory when $p$ is very large, but the dependency of the procedure on $K$ is poor. This motivates the idea of using MoM not as a final estimator, but as a warm start to an iterative method. In the following section, we show that the EM algorithm, initialized with the MoM estimator, has much better performance than the MoM estimator alone.
3 Parameter estimation via the EM algorithm
We formally introduce the parameter estimation via the EM algorithm under softmax mixtures. The EM algorithm is given in Section 3.1. Its theoretical guarantees are established in Section 3.2 and we outline its proof in Section 3.3.
3.1 Estimation of the softmax mixture parameters via EM
In this section we state the steps of the EM algorithm for soft-max mixtures. The log-likelihood of the sample $Y_{1},\ldots,Y_{N}$ drawn independently from $\pi_{{\bm{\omega}}}(\cdot\mid X_{1},\ldots,X_{p})$ equals
$\ell_{N}({\bm{\omega}})$ $=\sum_{i=1}^{N}\log\left(\sum_{k=1}^{K}\alpha_{k}A_{{\bm{\theta}}_{k}}(Y_{i}\mid X_{1},\ldots,X_{p})\right)$
$=N\sum_{i=1}^{p}\widehat{\pi}_{j}\log\left(\sum_{k=1}^{K}\alpha_{k}{\exp\left(X_{i}^{\top}{\bm{\theta}}_{k}\right)\over\sum_{j=1}^{p}\exp\left(X_{j}^{\top}{\bm{\theta}}_{k}\right)}\right),$
where $\widehat{\pi}_{j}$ denotes the observed frequency of $X_{j}$ in the sample.
Since this function is not concave in the parameter ${\bm{\omega}}=(\bm{\alpha},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$, it is hard to find the exact MLE solution and we propose the EM-algorithm to estimate the true model parameter
${\bm{\omega}}^{*}=(\bm{\alpha}^{*},{\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}).$
Since all results in this section are obtained given the observed $X_{1}=x_{1},\ldots,X_{p}=x_{p}$, we lighten our notation, by dropping the explicit conditioning on the observed values, whenever there is no possibility of confusion. For instance, we will write $A_{{\bm{\theta}}_{k}}(X_{j}):=A_{{\bm{\theta}}_{k}}(X_{j}\mid X_{1},\ldots,X_{p})$ for $j\in[p]$.
In order to state the EM-algorithm, we need to introduce additional quantities. First, let $Z$ be the latent random vector taking values within $\{{\bm{\theta}}_{1},\cdots,{\bm{\theta}}_{K}\}$ with corresponding probabilities in ${\bm{\alpha}}$. We define the conditional probability of $Z={\bm{\theta}}_{k}$ given that $Y=x$ as
$h_{{\bm{\omega}}}({\bm{\theta}}_{k}\mid x):={\alpha_{k}~{}A_{{\bm{\theta}}_{k}}(x)\over\pi_{\bm{\omega}}(x)}={\alpha_{k}~{}A_{{\bm{\theta}}_{k}}(x)\over\sum_{\ell=1}^{K}\alpha_{\ell}~{}A_{{\bm{\theta}}_{\ell}}(x)}.$ (38)
Second, we define the joint probability of $Z={\bm{\theta}}_{k}$ and $Y=x$ as
$\log f_{\bm{\omega}}(x,{\bm{\theta}}_{k})$ $:=~{}\log\mathbb{P}\{Y=x,Z={\bm{\theta}}_{k}\mid X_{1},\ldots,X_{p}\}$ (39)
$~{}=~{}\log(\alpha_{k})+\log\left(A_{{\bm{\theta}}_{k}}(x)\right)$
$~{}=~{}\log(\alpha_{k})+{\bm{\theta}}_{k}^{\top}x-\log\left(\sum_{\ell=1}^{p}\exp\left({X_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right).$
Instead of maximizing the marginal log-likelihood, the EM algorithm iteratively maximizes the surrogate function
$\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $~{}=\sum_{j=1}^{p}\widehat{\pi}_{j}\sum_{k=1}^{K}h_{{\bm{\omega}}^{\prime}}({\bm{\theta}}_{k}^{\prime}\mid X_{j})\log f_{\bm{\omega}}(X_{j},{\bm{\theta}}_{k})$ (40)
over its first argument ${\bm{\omega}}$. Plugging in (38) and (39) into (40), we can write this criterion as
$\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $=\sum_{j=1}^{p}\widehat{\pi}_{j}\sum_{k=1}^{K}{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}\left[\log(\alpha_{k})+X_{j}^{\top}{\bm{\theta}}_{k}-\log\left(\sum_{\ell=1}^{p}\exp\left({X_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right)\right].$
In the parlance of the EM-algorithm literature, evaluating the $\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ function corresponds to the “E-step”, while maximizing over ${\bm{\omega}}$ is the “M-step”. Starting at some initial point $\widehat{\bm{\omega}}^{(0)}$, the classical EM-algorithm iterates as follows:
$\widehat{\bm{\omega}}^{(t+1)}=\operatorname*{argmax}_{{\bm{\omega}}}~{}\widehat{Q}({\bm{\omega}}\mid\widehat{\bm{\omega}}^{(t)})\qquad\text{for}~{}t=0,1,2,\ldots$ (41)
until convergence.
For the problem at hand, maximization of $\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ in (40) over ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$ is a concave optimization problem. More specifically,
  • maximizing with respect to ${\bm{\alpha}}\in\Delta_{K}$ yields the following closed-form solution
    $\widehat{\alpha}_{k}^{(t+1)}$ $=\sum_{j=1}^{p}\widehat{\pi}_{j}{\widehat{\alpha}_{k}^{(t)}A_{\widehat{\bm{\theta}}_{k}^{(t)}}(X_{j})\over\pi_{\widehat{\bm{\omega}}^{(t)}}(X_{j})}:=\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)}),$ (42)
    for all $k\in[K]$.
  • maximization over vectors ${\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}$ does not admit a closed-form solution, and we propose a gradient-ascent step which is often used in such circumstances. For all $k\in[K]$, let $\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{(t)})$ be the gradient of $\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{(t)})$ with respect to ${\bm{\theta}}_{k}$ in the first argument ${\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$. Given a chosen step size $\eta_{k}>0$, the gradient-ascent step in our first-order-EM algorithm is
    $\widehat{\bm{\theta}}_{k}^{(t+1)}$ $=\widehat{\bm{\theta}}_{k}^{(t)}+\eta_{k}~{}\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid\widehat{\bm{\omega}}^{(t)})\mid_{{\bm{\omega}}=\widehat{\bm{\omega}}^{(t)}}.$ (43)
    The gradient equals
    $\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $=\sum_{j=1}^{p}\widehat{\pi}_{j}h_{{\bm{\omega}}^{\prime}}({\bm{\theta}}_{k}^{\prime}\mid X_{j})\left(X_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}\right)$
    $=\sum_{j=1}^{p}\widehat{\pi}_{j}{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}\left(X_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}\right).$ (44)
    Here, we write $A_{\bm{\theta}}=(A_{{\bm{\theta}}}(X_{1}),\ldots,A_{{\bm{\theta}}}(X_{p}))^{\top}\in\Delta_{p}$ whence
    $\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}=\sum_{j=1}^{p}X_{j}A_{{\bm{\theta}}_{k}}(X_{j})={\sum_{j=1}^{p}X_{j}\exp(X_{j}^{\top}{\bm{\theta}}_{k})\over\sum_{i=1}^{p}\exp(X_{i}^{\top}{\bm{\theta}}_{k})}.$
    Therefore, for all $k\in[K]$, the gradient-ascent step in (3.1) is concretely
    $\widehat{\bm{\theta}}_{k}^{(t+1)}$ $=\widehat{\bm{\theta}}_{k}^{(t)}+\eta_{k}~{}\sum_{j=1}^{p}\widehat{\pi}_{j}{\widehat{\alpha}_{k}^{(t)}A_{\widehat{\bm{\theta}}_{k}^{(t)}}(X_{j})\over\pi_{\widehat{\bm{\omega}}^{(t)}}(X_{j})}\left(X_{j}-\mathbf{X}^{\top}A_{\widehat{\bm{\theta}}_{k}^{(t)}}\right).$
Since the update (42) is given in closed form, while (3.1) is a gradient-ascent step, the EM algorithm for softmax mixtures can be viewed as a hybrid algorithm.
3.2 Theoretical guarantees for EM estimation of softmax mixture paramters
The parameter space is denoted by
$\Omega=\left\{{\bm{\omega}}=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K}):\bm{\alpha}\in\Delta_{K},~{}{\bm{\theta}}_{k}\in{\mathbb{B}}(\mathbf{0},B),~{}\forall k\in[K]\right\},$
where ${\mathbb{B}}(\mathbf{0},B)$ denotes the Euclidean ball around $\mathbf{0}$ with radius equal to $B$. Recall Assumptions 2, 3 and 6 from Section 2.3, which we also make in this section.
For any ${\bm{\omega}},{\bm{\omega}}^{\prime}\in\Omega$, we define the distance
$d({\bm{\omega}},{\bm{\omega}}^{\prime})=\max\left\{\overline{\sigma}\max_{k\in[K]}\|{\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime}\|_{2},~{}{1\over\underline{\alpha}}\|{\bm{\alpha}}-{\bm{\alpha}}^{\prime}\|_{\infty}\right\}.$ (45)
It is well known that the EM algorithm is very sensitive to its starting value $\widehat{\bm{\omega}}^{(0)}$. Our analysis reveals that the algorithm converges to ${\bm{\omega}}^{*}$ for any starting point that satisfies
$d(\widehat{\bm{\omega}}^{(0)},{\bm{\omega}}^{*})\leq\delta_{0}\quad\text{ with }\quad\delta_{0}\sqrt{L+\log(p)}\leq c_{0}$ (46)
for some sufficiently small constant $20<c_{0}<1/$ 2.
Propositions 2 and 2 of the previous section shows that the MoM estimator of the softmax mixture parameters is an excellent warm start candidate: it trivially meets (46) for any $L$ that satisfies Assumption 8. Now we are ready to present our main result on the rate of convergence of the EM-updates.
Theorem 1.
Grant Assumptions 2,3,6,7,8. Assume the initialization $\widehat{\bm{\omega}}^{(0)}$ satisfies (46) with initial bound $\delta_{0}$. Then, there exist some $10<\phi<$ 1, some absolute constant $C>0$ and step-sizes $\eta_{k}>0$, $k\in[K]$, such that for the sequence $\widehat{\bm{\omega}}^{(t)}$ computed via (42) and (3.1), we have
$\mathbb{P}\left[\bigcap_{t\geq 0}\left.\left\{d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq~{}\phi^{t}\delta_{0}+C\sqrt{L\log N\over N}\right\}~{}\right|~{}X_{1},\ldots,X_{p}\right]\geq 1-\mathcal{O}(N^{-L})$
$\mu$-almost surely, as $p\to\infty$.
Our main result states that under this set of conditions, the estimates $\widehat{\bm{\omega}}^{(t)}$, using any initial estimate $\widehat{\bm{\omega}}^{(0)}$ with (46) and with steps (42) and (3.1) converge at the rate $\sqrt{L\log(N)/N}$ after (in order) $\log(N/L^{2})$ number of iterations. This rate is only off by the logarithmic factor $\log(N)$ from the parametric rate. In particular, this rate holds for each of $\overline{\sigma}\max_{k\in[K]}\|\widehat{\bm{\theta}}_{k}^{(t)}-{\bm{\theta}}_{k}\|_{2}$ and $\|\widehat{\bm{\alpha}}^{(t)}-{\bm{\alpha}}\|_{\infty}/\underline{\alpha}$ individually, after the same $\log(N/L^{2})$ number of iterations.
Our theory shows that the step size $\eta_{k}$ cannot be chosen too large in order to ensure convergence of the EM updates. On the other hand, a smaller choice of $\eta_{k}$ will lead to slower algorithmic convergence (in the sense that $\phi$ gets closer to 1), but will not affect the final statistical accuracy. We found that $\eta_{k}=1$ yields overall satisfactory results in our numerical experiments.
3.3 Outline of the proof
At the heart of the analysis of the EM-algorithm lies a simple, well-known principle (which we present here with a two-line proof).
Proposition 7.
Let $(\Omega,d)$ be a metric space. Let $\{{\bm{\omega}}^{(t)}\}_{t\geq 0}$ be any sequence in $\Omega$ that satisfies $d({\bm{\omega}}^{(0)},{\bm{\omega}}^{*})\leq\delta_{0}$ for some $\delta_{0}>0$ and, for all $t\geq 0$,
$d({\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})$ $\leq\delta_{N}+\phi~{}d({\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$ (47)
for some $\delta_{N}\leq(1-\phi)\delta_{0}$ and $10<\phi<$ 1. Then, $d({\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})\leq\delta_{0}$ and
$d({\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})$ $\leq\phi^{t+1}\delta_{0}+{1-\phi^{t+1}\over 1-\phi}\delta_{N}.$ (48)
@@ 3.3 Outline of the proof
The proof uses an induction argument. Assume ${\bm{\omega}}^{(t)}\in\mathbb{B}_{d}({\bm{\omega}}^{*},\delta_{0})$, the ball centered around ${\bm{\omega}}^{*}$, with radius $\delta_{0}$, with respect to the metric $d$. Then
$d({\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})$ $\leq\delta_{N}+\phi~{}d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})\leq(1-\phi)\delta_{0}+\phi\delta_{0}=\delta_{0}.$
If (48) holds at step $t$ then, by property (47), it also holds at step $t+1$. ∎
This basic observation shows that it is key to establish (47) with $\phi<1$. The problem at hand is non-standard in that we are dealing with a hybrid between the standard EM for ${\bm{\alpha}}$ in step (42) and the first-order EM for ${\bm{\theta}}_{k}$, $k\in[K]$, in step (3.1). Pick any $k\in[K]$. First we study the updates $\widehat{\alpha}_{k}^{(t+1)}$. Let
$M_{k}({\bm{\omega}})$ $=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}h_{{\bm{\omega}}}({\bm{\theta}}_{k}\mid X_{j})=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j}){\alpha_{k}A_{{\bm{\theta}}_{k}}(X_{j})\over\pi_{\bm{\omega}}(X_{j})}$ (49)
be the population counterpart of $\widehat{M}_{k}({\bm{\omega}})$ in (42). Notice that $M_{k}({\bm{\omega}}^{*})=\alpha_{k}^{*}$ (self-consistency), so we can write
$|\widehat{\alpha}_{k}^{(t+1)}-\alpha_{k}^{*}|$ $=|\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$
$\leq|\widehat{M}_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}(\widehat{\bm{\omega}}^{(t)})|+|M_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$
$\leq\sup_{d({\bm{\omega}},{\bm{\omega}}^{*})\leq\delta_{0}}|\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})|+|M_{k}(\widehat{\bm{\omega}}^{(t)})-M_{k}({\bm{\omega}}^{*})|$
The nontrivial Lemma 4 gives a rate for the empirical process $\widehat{M}_{k}({\bm{\omega}})-M_{k}({\bm{\omega}})$ over the ball $\mathbb{B}_{d}({\bm{\omega}}^{*},\delta_{0})$, while Lemma 3 establishes the Lipschitz property for $M_{k}(\cdot)$ with high probability. Assumption 7 ensures that the Lipschitz constant is strictly less than 1.
The analysis for the first-order EM-updates $\widehat{\bm{\theta}}_{k}^{(t+1)}$ is slightly more involved. Let
$Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ $=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}\sum_{k=1}^{K}h_{{\bm{\omega}}^{\prime}}({\bm{\theta}}_{k}^{\prime}\mid X_{j})\log f_{\bm{\omega}}(X_{j},{\bm{\theta}}_{k})$ (50)
$=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}\sum_{k=1}^{K}{\alpha_{k}^{\prime}A_{{\bm{\theta}}^{\prime}_{k}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}\left[\log(\alpha_{k})+X_{j}^{\top}{\bm{\theta}}_{k}-\log\left(\sum_{\ell=1}^{p}\exp\left({X_{\ell}^{\top}{\bm{\theta}}_{k}}\right)\right)\right].$
be the population level counterpart of $\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ in (40). Its gradient with respect to ${\bm{\theta}}_{k}$ equals
$\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})=\sum_{j=1}^{p}\pi_{{\bm{\omega}}^{*}}(X_{j})~{}{\alpha_{k}^{\prime}A_{{\bm{\theta}}_{k}^{\prime}}(X_{j})\over\pi_{{\bm{\omega}}^{\prime}}(X_{j})}(X_{j}-\mathbf{X}^{\top}A_{{\bm{\theta}}_{k}}).$ (51)
By the definition of the sequence $\widehat{\bm{\theta}}^{(t)}_{k}$, $t\geq 0$, and the triangle inequality, we find
$\leq\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid{\bm{\omega}}^{*})\|_{2}$
$\qquad+\eta_{k}\|\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid\widehat{\bm{\omega}}^{(t)})-\nabla_{{\bm{\theta}}_{k}}Q(\widehat{\bm{\omega}}^{(t)}\mid{\bm{\omega}}^{*})\|_{2}$
Now, we set $q_{k}({\bm{\omega}}):=\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{*})$. This function satisfies $q_{k}({\bm{\omega}}^{*})=0$, and its smoothness and strong-concavity properties are stated in Lemma 2. This takes care of the first term
$\|\widehat{\bm{\theta}}^{(t)}_{k}-{\bm{\theta}}_{k}^{*}+\eta_{k}q(\widehat{\bm{\omega}}^{(t)})\|_{2}$
on the right of the display above. Finally, the non-trivial Lemma 3, that establishes a Lipschitz property of the map ${\bm{\omega}}^{\prime}\mapsto\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}}^{\prime})$ in the $\ell_{2}$-norm, and the second maximal inequality for
$\sup_{d({\bm{\omega}},{\bm{\omega}}^{*})\leq\delta_{0}}\|\nabla_{{\bm{\theta}}_{k}}\widehat{Q}({\bm{\omega}}\mid{\bm{\omega}})-\nabla_{{\bm{\theta}}_{k}}Q({\bm{\omega}}\mid{\bm{\omega}})\|_{2}$
in Lemma 4, are used to control the last two terms on the right. A careful analysis is needed to warrant that there exists $\phi=\phi(c_{0},\underline{\sigma},C_{\Delta})<1$ such that
$d(\widehat{\bm{\omega}}^{(t+1)},{\bm{\omega}}^{*})$ $\leq C\sqrt{L\log N\over N}+\phi~{}d(\widehat{\bm{\omega}}^{(t)},{\bm{\omega}}^{*})$
with high probability. We provide a full proof in Appendix B.

### v2

@@ 3.4.1 Application to practical choice of the primary axis
so that the selected $v_{i^{*}}$ approximately maximizes $\det({\mathbf{M}}(v_{i}))$, thereby leading to the largest separation among $v_{i}^{\top}{\bm{\theta}}_{1}^{*},\ldots,v_{i}^{\top}{\bm{\theta}}_{K}^{*}.$
3.4.2 Application to random initialization of the EM
In Theorem 2 of Section 2.2 we show that the EM algorithm has provable guarantees once its initialization meets (20). In addition to using the MoM estimator developed in Section 3.3, it is common in practice to simply deploy random initializations, that is, by simply drawing ${\bm{\theta}}_{1}^{(0)},\ldots,{\bm{\theta}}_{K}^{(0)}$ uniformly from a chosen sphere multiple times, and selecting the corresponding EM estimate that yields the highest likelihood.
The intuition is the following: Let ${\bm{\theta}}^{*}\in\mathbb{S}^{L-1}$ be a given target vector and let $\delta>0$ be the desired accuracy. Then for any $\varepsilon>0$, with probability at least $1-\epsilon$, there exists at least one vector $\bar{v}$ in independent draws $\{v_{1},\ldots,v_{m}\}$ from $\mathbb{S}^{L-1}$ such that
$\|\bar{v}-{\bm{\theta}}^{*}\|_{2}\leq\delta$ (61)
provided that
$m\geq\exp\left(L(1-{\delta^{2}/2})\right)\log(1/\varepsilon).$
The above result follows from a simple union bound argument together with the spherical cap probability bound in Tkocz (2012). For completeness, we include its proof in Section C.12. As a result, in the worst case one needs to use $\exp(\mathcal{O}(L))$ random initializations and run the EM algorithm this many times, which is computationally expensive when $L$ is not small.
However, if ${\bm{\theta}}^{*}$ is known to lie within a subspace of dimension at most $K\ll L$, then one only needs $\exp(\mathcal{O}(K))$ random draws from the unit sphere in this subspace to achieve the desired $\delta_{0}$ accuracy. We formalize this in the following lemma in our context. Recall that ${\bm{\theta}}_{1}^{*},\ldots,{\bm{\theta}}_{K}^{*}$ lie in the column space of $\Gamma$ given in (52). Further recall that $\widehat{V}\in\mathbb{O}_{L\times K}$ contains the first $K$ leading eigenvectors of $\widehat{\Gamma}$, the estimator of $\Gamma$ given in (54).
Lemma 3.
Fix any $k\in[K]$ and ${\bm{\theta}}_{k}^{*}\in\mathbb{S}^{L-1}$. Let $v_{1},\ldots,v_{m}$ be independently sampled as (60). For arbitrary $\varepsilon>0$ and $\delta_{0}>0$, on the event $\{\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}\leq(\underline{\alpha}/2)\delta_{\Gamma}\}$ for some $\delta_{\Gamma}<\delta_{0}^{2}/2$, with probability at least $1-\varepsilon$, there exists at least one vector $\bar{v}\in\{v_{1},\ldots,v_{m}\}$ such that $\|\bar{v}-{\bm{\theta}}_{k}^{*}\|_{2}\leq\delta_{0}$ provided that
$m\geq\exp\left\{K(1-\delta_{0}^{2}/2+\delta_{\Gamma})\right\}\log(1/\varepsilon).$
@@ 3.4.2 Application to random initialization of the EM
The proof is stated in Section C.12. ∎
By plugging into the bound of $\delta_{0}$ in (20) as well as the bound of $\|\widehat{\Gamma}-\Gamma\|_{\mathrm{op}}$ in Proposition 3, the requirement $\delta_{\Gamma}<\delta_{0}^{2}/2$ becomes $\|\mathbf{X}\|_{\infty,2}^{3}\sqrt{\log N/N}+\|\mathbf{X}\|_{\infty,2}^{2}\epsilon_{p}^{\prime}\leq c(\underline{\alpha},\varsigma,\overline{\sigma})$ where we further recall that $\epsilon_{p}^{\prime}=\mathcal{O}(\sqrt{\log\log p/p})$ and $\epsilon_{p}^{\prime}=\mathcal{O}(\sqrt{L\log p/p})$ in the settings of Theorem 3 and 4, respectively.